# Reappraisal of SU(3)-flavor breaking in $B\rightarrow DP$

### v1

@@ Abstract
In light of recently found deviations of the experimental data from predictions from QCD factorization for $B_{(s)}\rightarrow D_{(s)}P$ decays, where $P=\{\pi,K\}$, we systematically probe the current status of the SU(3)${}_{F}$ expansion from a fit to experimental branching ratio data without any further theory input. We find that the current data is in agreement with the power-counting of the SU(3)${}_{F}$ expansion. While the SU(3)${}_{F}$ limit is excluded at $>5\sigma$, SU(3)${}_{F}$-breaking contributions of $\sim 20\%$ suffice for an excellent description of the data. SU(3)${}_{F}$ breaking is needed in tree ($>5\sigma$) and color-suppressed tree ($2.4\sigma$) diagrams. We are not yet sensitive to SU(3)${}_{F}$ breaking in exchange diagrams. From the underlying SU(3)${}_{F}$ parametrization we predict the unmeasured branching ratios $\mathcal{B}(\overline{B}_{s}^{0}\rightarrow\pi^{-}D^{+})=2\mathcal{B}(\overline{B}_{s}^{0}\rightarrow\pi^{0}D^{0})=[0.3,7.2]\times 10^{-6}$ of suppressed decays that can be searched for at the LHCb Upgrade I and LHCb Upgrade II experiments respectively.
@@ I Introduction
Theoretical predictions for the branching ratios of $B_{(s)}\rightarrow D_{(s)}P$, $P=\{\pi,K\}$ and $B_{(s)}\rightarrow D_{(s)}V$, $V=\{\rho,K^{*}\}$ decays based on QCD factorization (QCDF) show a significant disagreement with the experimental data Bordone et al. (2020); Cai et al. (2021). This resulted in a lot of renewed interest in these decays, both theoretically Gershon et al. (2022); Piscopo and Rusov (2023); Lenz et al. (2023); Iguro and Kitahara (2020); Endo et al. (2022); Fleischer and Malami (2022); Beneke et al. (2021); Bordone et al. (2021) and experimentally Krohn et al. (2023); Aaij et al. (2023); Dib et al. (2023); Waheed et al. (2022); Aaij et al. (2021a). Also in other hadronic $B$ decays anomalies have been seen Berthiaume et al. (2023); Bhattacharya et al. (2023); Amhis et al. (2023a); Biswas et al. (2023); Brod et al. (2015); Jäger et al. (2018); Lenz and Tetlalmatzi-Xolocotzi (2020). In the case of the theoretically clean $\overline{B}_{s}^{0}\rightarrow\pi^{-}D_{s}^{+}$ and $\overline{B}^{0}\rightarrow K^{-}D^{+}$ decays, which are free from penguin and annihilation topologies and dominated by colour-allowed tree processes, this disagreement is sizable, see the comparison of theoretical and experimental results in Table 1. It is important to note that theoretical predictions based on QCDF for the ratios of the branching fractions,
@@
Decays involving $b\rightarrow c\overline{u}q$ weak transitions are used extensively for measurements of CP-violation in the Standard Model (SM). Within the SM the CKM phase $\gamma$ can be measured cleanly via the interference between the amplitudes of $b\rightarrow c\overline{u}s$ and $b\rightarrow u\overline{c}s$ weak transitions with a theoretical uncertainty of $\lesssim 10^{-7}$ Brod and Zupan (2014); the methodologies to do so include the GLW Gronau and London (1991); Gronau and Wyler (1991), ADS Atwood et al. (1997, 2001) and BPGGSZ method Bondar (2002); Giri et al. (2003); Poluektov et al. (2004); Bondar and Poluektov (2006); Ceccucci et al. (2020), see also the overview in Sec. 77 of Ref. Workman et al. (2022). Recent developments regarding unbinned extractions of $\gamma$ can be found in Refs. Poluektov (2018); Backus et al. (2023); Lane et al. (2023).
@@
$B\rightarrow DP$ decays have been studied for a long time. The sensitivity to mixing phases and $\gamma$ have been further explored in Refs. Dunietz (1998); Fleischer (2003a, b). Implications of rescattering effects were looked at in Refs. Chua and Hou (2008); Cheng et al. (2005); Chua et al. (2002); Endo et al. (2022). QCD factorization and SCET have been applied in Refs. Beneke et al. (2000); Bauer et al. (2001); Keum et al. (2004); Huber et al. (2016); Cai et al. (2021); Bordone et al. (2020); Mantry et al. (2003), and the factorization-assisted topological (FAT) approach in Ref. Zhou et al. (2015). Flavor-symmetry based methods have been used in Refs. Zeppenfeld (1981); Savage and Wise (1989); Gronau et al. (1995); Chiang and Senaha (2007); Jung and Mannel (2009); Fleischer et al. (2011); Kenzie et al. (2016); Kitazawa et al. (2019); Colangelo and Ferrandes (2005); Grinstein and Lebed (1996); Neubert and Petrov (2001); Xing (2002); Cheng (2002); Wolfenstein (2004); Chiang and Rosner (2003); Xing (2003); Yamamoto (1994); Xing (1995); Wolfenstein (2004); Kim et al. (2005). The most recent fits in plain SU(3)${}_{F}$ have been performed in Refs. Chiang and Senaha (2007); Colangelo and Ferrandes (2005); Colangelo et al. (2007). Ref. Bordone et al. (2020) also looks at SU(3)${}_{F}$ in the context of QCDF. New physics sensitivities have been recently explored in Ref. Iguro and Kitahara (2020).
In this paper, in light of the deviations of branching ratio data from QCDF, we analyze systematically the quality of the SU(3)${}_{F}$ expansion in $B_{(s)}\rightarrow D_{(s)}P$ decays, with a methodology similar to the one used in Ref. Müller et al. (2015) for charm decays. We employ the SU(3)${}_{F}$ expansion derived in Ref. Gronau et al. (1995), and include SU(3)${}_{F}$-breaking effects purely through topological diagrams that we extract from the data. In Sec. II we introduce our notation, recapitulate the SU(3)${}_{F}$ decomposition and introduce the considered measures of SU(3)${}_{F}$ breaking. We show our numerical results in Sec. III and conclude in Sec. IV
@@
II Notation and measures of SU(3)${}_{F}$ breaking
@@ II.1 Parametrizations
Table 2: SU(3)${}_{F}$ decomposition. Table adapted from Ref. Gronau et al. (1995).
We use the topological SU(3)${}_{F}$ decomposition given in Ref. Gronau et al. (1995) which we summarize for convenience in Table 2. For our fit of the parametrization to the data, we use the following combination of parameters:
@@
II.2 Measures of SU(3)${}_{F}$ breaking
As parametrization-dependent measures of SU(3)${}_{F}$ breaking we use the parameters of Eq. (6)
@@ II.2 Measures of SU(3)${}_{F}$ breaking
As an alternative, parametrization-independent measure of SU(3)${}_{F}$ breaking we also consider the known SU(3)${}_{F}$ sum rules Gronau et al. (1995) between two decay channels and calculate the splitting around their average value
@@
$s_{4}$ $\equiv\frac{|\mathcal{A}(\bar{B}^{0}\rightarrow\pi^{-}D^{+})|/|\lambda_{d}|}{\mathcal{A}(\bar{B}_{s}^{0}\rightarrow K^{-}D_{s}^{+})|/|\lambda_{s}|}\,,$ (26)
@@
where in the SU(3)${}_{F}$ limit (including isospin)
@@ III Numerical Results
Table 6: Numerical results for the parametrization-independent measures of SU(3)${}_{F}$ breaking, resulting from $\Delta\chi^{2}$ scans. We do not show results for $\varepsilon_{1,2}$ as these contain yet unmeasured branching ratios, see Eqs. (23) and (24).
@@
Figure 1: $\Delta\chi^{2}$ for varying degrees of allowed SU(3)${}_{F}$ breaking $\delta_{X}$. $\Delta\chi^{2}=1\,(25)$i.e. $1\,(5)\sigma$, corresponds to $\delta_{X}=0.17(0.09)$, indicated by the dashed(dotted) line, and $\Delta\chi^{2}<1\times 10^{-3}$ is reached for $\delta_{X}>0.2$.
The current experimental status of the relevant branching ratio measurements is shown in Table 3. For $B_{s}$ decays, due to $\Delta\Gamma_{s}\neq 0$, we have to take into account a correction factor for the relation between the “theoretical” branching ratio at $t=0$ and the “experimental” time-integrated branching ratio De Bruyn et al. (2012). The implications of $\Delta\Gamma_{s}\neq 0$ for $B_{s}\rightarrow D_{s}^{\pm}K^{\mp}$ have been discussed in detail in Ref. Fleischer and Malami (2022) and we use the “theoretical” $\mathcal{B}(\overline{B}_{s}^{0}\rightarrow K^{-}D_{s}^{+})$ as extracted from the experimental data therein. This amounts to a correction of $\sim 14\%$ vs. the “experimental” value $\mathcal{B}(\overline{B}_{s}^{0}\rightarrow K^{\mp}D_{s}^{\pm})=(2.25\pm 0.12)\cdot 10^{-4}$ Workman et al. (2022), see Refs. Fleischer and Malami (2022); De Bruyn et al. (2013) for details.
The decay $\bar{B}_{s}^{0}\rightarrow\pi^{-}D_{s}^{+}$ is flavor-specific, see Ref. Gershon et al. (2022) for a detailed discussion. Therefore, in this case the correction factor accounting for the non-vanishing $B_{s}$ width difference $\Delta\Gamma_{s}$ reads $1-y_{s}^{2}$ De Bruyn et al. (2012), with Amhis et al. (2023b)
@@
and the average $B_{s}$ width $\Gamma_{s}$. As $y_{s}^{2}\sim 0.004$, and the relative uncertainty of $\bar{B}_{s}^{0}\rightarrow\pi^{-}D_{s}^{+}$ is an order of magnitude larger, we neglect the effect of the non-vanishing width difference for this decay.
@@
where $\delta_{X}$ is a measure of the allowed SU(3)${}_{F}$ breaking which we set to
@@
We have altogether ten experimental data points (two branching ratios not yet being measured) and 17 real theory parameters, of which nine are magnitudes and eight are strong phases. However, we also have six constraints Eqs. (32)–(34) on these parameters and it is indeed non-trivial if the data is in agreement with these constraints.
At the global minimum, for our null hypothesis we find $\chi^{2}=0$, i.e. a perfect fit to the data, meaning that the data can be explained with SU(3)${}_{F}$ breaking of $\delta_{X}=30\%$, as we expect. In order to demonstrate the non-triviality of this result, we show in Fig. 1 the resulting $\Delta\chi^{2}$ profile as a function of $\delta_{X}$, compared to our null hypothesis.
Our fit results for $\delta_{X}=0.3$ are presented in Tables 4, 5, and 6. All errors are obtained by profiled-$\chi^{2}$ scans. Since the SU(3)${}_{F}$ system is under-constrained, with more parameters than observables, it has a high degree of degeneracy. The parameter scans show non-Gaussianity, and appear flat around the minimum while sharply increasing near parameter boundaries. For that reason we quote parameter ranges instead of a central point with errors.
@@
  • The SU(3)${}_{F}$ limit is excluded at $>5\sigma$, mainly triggered by the need for SU(3)${}_{F}$ breaking in the tree diagrams, which is also $>5\sigma$. SU(3)${}_{F}$ breaking in the $C$ diagrams is needed at $2.4\sigma$, while we are not yet sensitive to SU(3)${}_{F}$ breaking in $E$ diagrams.
  • Compared to Ref. Colangelo and Ferrandes (2005), where the topological amplitudes are extracted in the SU(3)${}_{F}$ limit, and to Ref. Chiang and Senaha (2007), where partial SU(3)${}_{F}$ breaking effects are taken into account, we find that the general features of the results for $|C/T|$ and $|E/T|$ obtained in Colangelo and Ferrandes (2005); Chiang and Senaha (2007) are confirmed when including the first order corrections in full generality, however, the additional parameters result in larger errors for $|C/T|$.
  • The data can be described perfectly well with SU(3)${}_{F}$ breaking of $\sim 20\%$. This conclusion holds independently of employing parameter-dependent or parameter-independent measures of SU(3)${}_{F}$ breaking, see Fig. 1 and Table 6. Our observation agrees with the conclusions of studies of SU(3)${}_{F}$ breaking based on factorization Bordone et al. (2020); Fleischer et al. (2011), where the factorizable SU(3)${}_{F}$ breaking from the decay constants can be estimated as $f_{K}/f_{\pi}-1\sim 20\%$ Aoki et al. (2022), see also Ref. Jung and Mannel (2009).
  • The SU(3)${}_{F}$-breaking effects in $T$, $C$ and $E$ do not add up constructively in order to produce larger SU(3)${}_{F}$-breaking effects. This can be seen from the SU(3)${}_{F}$-breaking measure $\varepsilon_{4}$, which relates the $U$-spin pair $\overline{B}^{0}\to\pi^{-}D^{+}$ and $\overline{B_{s}}^{0}\to K^{-}D_{s}^{+}$, see Table 6.
@@
  • The data implies that the ratio $|E/T|$ is more strongly suppressed than $|C/T|$. However, $|C/T|$ is rather large compared to the expectation from $1/N_{c}$ counting Gronau et al. (1995). The obtained knowledge of $|E/T|$ is mainly driven by the constraint from
@@
  • The phases $\phi_{C}$ and $\phi_{E}$ are sizable, indicating large rescattering effects. The phases of the SU(3)${}_{F}$-breaking topologies are not yet very well known.
  • Due to lack of data we can not yet test the isospin relation $s_{1}=1$, see Eqs. (23, 29). From our fit we obtain the corresponding yet unmeasured branching ratios as
@@
    The interfering $B_{s}$ decays into the same final states are in this case of the same order in Wolfenstein-$\lambda$, like it is also the case for $\overline{B}_{s}^{0}\rightarrow K^{\mp}D_{s}^{\pm}$ decays, see above. Estimating the impact from $\Delta\Gamma_{s}\neq 0$ as $\sim 20\%$, as suggested by the size of the effect for $\mathcal{B}(\overline{B}_{s}^{0}\rightarrow K^{-}D_{s}^{+})$, we predict
@@
    Future measurements of $\mathcal{B}(\overline{B}^{0}_{s}\to\pi^{-}D^{+})$ and $\mathcal{B}(\overline{B}^{0}_{s}\to\pi^{0}D^{0})$ would allow to test their isospin symmetry relation and at the same time improve the bounds on $|E/T|$.

### v2

@@ Abstract
In light of recently found deviations of the experimental data from predictions from QCD factorization for $B_{(s)}\rightarrow D_{(s)}P$ decays, where $P=\{\pi,K\}$, we systematically probe the current status of the SU(3)F expansion from a fit to experimental branching ratio data without any further theory input. We find that the current data are in agreement with the power counting of the SU(3)F expansion. While the SU(3)F limit is excluded at $>5\sigma$, amplitude-level SU(3)F-breaking contributions of $\sim 20\%$ suffice for an excellent description of the data. SU(3)F breaking is needed in tree ($>5\sigma$) and color-suppressed tree ($2.4\sigma$) diagrams. We are not yet sensitive to SU(3)F breaking in exchange diagrams. From the underlying SU(3)F parametrization we predict the unmeasured branching ratios $\mathcal{B}(\overline{B}_{s}^{0}\rightarrow\pi^{-}D^{+})=2\mathcal{B}(\overline{B}_{s}^{0}\rightarrow\pi^{0}D^{0})=[0.3,7.2]\times 10^{-6}$ of suppressed decays that can be searched for at the LHCb experiment.
@@ I Introduction
Theoretical predictions for the branching ratios of $B_{(s)}\rightarrow D_{(s)}P$, $P=\{\pi,K\}$ and $B_{(s)}\rightarrow D_{(s)}V$, $V=\{\rho,K^{*}\}$ decays based on QCD factorization (QCDF) show a significant disagreement with the experimental data Bordone et al. (2020); Cai et al. (2021). This resulted in a lot of renewed interest in these decays, both theoretically Gershon et al. (2022); Piscopo and Rusov (2023); Lenz et al. (2023); Iguro and Kitahara (2020); Endo et al. (2022); Fleischer and Malami (2022); Beneke et al. (2021); Bordone et al. (2021) and experimentally Krohn et al. (2023); Aaij et al. (2023); Dib et al. (2023); Waheed et al. (2022); Aaij et al. (2021a). Also in other hadronic $B$ decays anomalies have been seen Berthiaume et al. (2023); Bhattacharya et al. (2023); Amhis et al. (2023a); Biswas et al. (2023); Brod et al. (2015); Jäger et al. (2018); Lenz and Tetlalmatzi-Xolocotzi (2020). In the case of the theoretically clean $\overline{B}_{s}^{0}\rightarrow\pi^{-}D_{s}^{+}$ and $\overline{B}^{0}\rightarrow K^{-}D^{+}$ decays, which are free from penguin and annihilation topologies and dominated by color-allowed tree processes, this disagreement is sizable; see the comparison of theoretical and experimental results in Table 1. It is important to note that theoretical predictions based on QCDF for the ratios of the branching fractions,
@@
Decays involving $b\rightarrow c\overline{u}q$ weak transitions are used extensively for measurements of CP violation in the Standard Model (SM). Within the SM the Cabibbo Kobayashi Maskawa (CKM) phase $\gamma$ can be measured cleanly via the interference between the amplitudes of $b\rightarrow c\overline{u}s$ and $b\rightarrow u\overline{c}s$ weak transitions with a theoretical uncertainty of $\lesssim 10^{-7}$ Brod and Zupan (2014); the methodologies to do so include the Gronau London Wyler Gronau and London (1991); Gronau and Wyler (1991), Atwood Dunietz Soni Atwood et al. (1997, 2001), and Bondar Poluektov Giri Grossman Soffer Zupan method Bondar (2002); Giri et al. (2003); Poluektov et al. (2004); Bondar and Poluektov (2006); Ceccucci et al. (2020), see also the overview in Sec. 77 of Ref. Workman et al. (2022). Recent developments regarding unbinned extractions of $\gamma$ can be found in Refs. Poluektov (2018); Backus et al. (2023); Lane et al. (2023).
@@
$B\rightarrow DP$ decays have been studied for a long time. The sensitivity to mixing phases and $\gamma$ have been further explored in Refs. Dunietz (1998); Fleischer (2003a, b). Implications of rescattering effects were looked at in Refs. Chua and Hou (2008); Cheng et al. (2005); Chua et al. (2002); Endo et al. (2022). QCD factorization and soft-collinear effective theory have been applied in Refs. Beneke et al. (2000); Bauer et al. (2001); Keum et al. (2004); Huber et al. (2016); Cai et al. (2021); Bordone et al. (2020); Mantry et al. (2003), and the factorization-assisted topological (FAT) approach in Ref. Zhou et al. (2015). Flavor-symmetry based methods have been used in Refs. Zeppenfeld (1981); Savage and Wise (1989); Gronau et al. (1995); Chiang and Senaha (2007); Jung and Mannel (2009); Fleischer et al. (2011); Kenzie et al. (2016); Kitazawa et al. (2019); Colangelo and Ferrandes (2005); Grinstein and Lebed (1996); Neubert and Petrov (2001); Xing (2002); Cheng (2002); Wolfenstein (2004); Chiang and Rosner (2003); Xing (2003); Yamamoto (1994); Xing (1995); Wolfenstein (2004); Kim et al. (2005). The most recent fits in plain SU(3)F have been performed in Refs. Chiang and Senaha (2007); Colangelo and Ferrandes (2005); Colangelo et al. (2007). Reference Bordone et al. (2020) also looks at SU(3)F in the context of QCDF. New physics sensitivities have recently been explored in Ref. Iguro and Kitahara (2020).
In this paper, in light of the deviations of branching ratio data from QCDF, we analyze systematically the quality of the SU(3)F expansion in $B_{(s)}\rightarrow D_{(s)}P$ decays, with a methodology similar to the one used in Ref. Müller et al. (2015) for charm decays. We employ the SU(3)F expansion derived in Ref. Gronau et al. (1995), and include SU(3)F-breaking effects purely through topological diagrams that we extract from the data. In Sec. II we introduce our notation, recapitulate the SU(3)F decomposition and introduce the considered measures of SU(3)F breaking. We show our numerical results in Sec. III and conclude in Sec. IV
@@
II Notation and measures of SU(3)F breaking
@@ II.1 Parametrizations
Table 2: SU(3)F decomposition. The table is adapted from Ref. Gronau et al. (1995).
We use the topological SU(3)F decomposition given in Ref. Gronau et al. (1995) which we summarize for convenience in Table 2. For our fit of the parametrization to the data, we use the following combination of parameters:
@@
SU(3)F is an approximate symmetry of the QCD Lagrangian that becomes exact in the limit $m_{u}=m_{d}=m_{s}$. Effects from corrections to this limit that account for $m_{s}\neq m_{d}$ can be incorporated in a systematic fashion with perturbation theory in terms of mass insertions of the form $\sim(m_{s}-m_{d})\bar{s}s$ Gronau et al. (1995); Müller et al. (2015); Jung and Mannel (2009). The diagrams $T_{i}$, $C_{i}$, and $E_{i}$ have the same flavor and color flow as the corresponding diagrams $T$, $C$, and $E$, except that they each include one mass insertion on one of the strange quark fermion lines (see Fig. 2 in Ref. Gronau et al. (1995)). In this study we include effects at first order in the expansion in SU(3)F breaking only. Second order corrections would come from diagrams that include two strange quark mass insertions and are beyond the scope of this study.
@@
II.2 Measures of SU(3)F breaking
As parametrization-dependent measures of SU(3)F breaking we use the parameters of Eq. (6)
@@ II.2 Measures of SU(3)F breaking
As an alternative, parametrization-independent measure of SU(3)F breaking we also consider the known SU(3)F sum rules Gronau et al. (1995) between two decay channels and calculate the splitting around their average value
@@
$s_{4}$ $\equiv\frac{|\mathcal{A}(\bar{B}^{0}\rightarrow\pi^{-}D^{+})|/|\lambda_{d}|}{|\mathcal{A}(\bar{B}_{s}^{0}\rightarrow K^{-}D_{s}^{+})|/|\lambda_{s}|}\,,$ (26)
@@
where in the SU(3)F limit (including isospin)
@@ III Numerical Results
Table 6: Numerical results for the parametrization-independent measures of SU(3)F breaking, resulting from $\Delta\chi^{2}$ scans. We do not show results for $\varepsilon_{1,2}$ as these contain yet unmeasured branching ratios, see Eqs. (23) and (24).
@@
Figure 1: $\Delta\chi^{2}$ for varying degrees of allowed SU(3)F breaking $\delta_{X}$. $\Delta\chi^{2}=1\,(25)$i.e.$1\,(5)\sigma$, corresponds to $\delta_{X}=0.17(0.09)$, indicated by the dashed(dotted) line, and $\Delta\chi^{2}<1\times 10^{-3}$ is reached for $\delta_{X}>0.2$.
The current experimental status of the relevant branching ratio measurements is shown in Table 3. For $B_{s}$ decays, due to $\Delta\Gamma_{s}\neq 0$, we have to take into account a correction factor for the relation between the “theoretical” branching ratio at $t=0$ and the “experimental” time-integrated branching ratio De Bruyn et al. (2012). The implications of $\Delta\Gamma_{s}\neq 0$ for $B_{s}\rightarrow D_{s}^{\pm}K^{\mp}$ have been discussed in detail in Ref. Fleischer and Malami (2022) and we use the “theoretical” $\mathcal{B}(\overline{B}_{s}^{0}\rightarrow K^{-}D_{s}^{+})$ as extracted from the experimental data therein. This amounts to a correction of $\sim 14\%$ vs. the “experimental” value $\mathcal{B}(\overline{B}_{s}^{0}\rightarrow K^{\mp}D_{s}^{\pm})=(2.25\pm 0.12)\cdot 10^{-4}$ Workman et al. (2022); see Refs. Fleischer and Malami (2022); De Bruyn et al. (2013) for details.
The decay $\bar{B}_{s}^{0}\rightarrow\pi^{-}D_{s}^{+}$ is flavor specific (see Ref. Gershon et al. (2022) for a detailed discussion). Therefore, in this case the correction factor accounting for the nonvanishing $B_{s}$ width difference $\Delta\Gamma_{s}$ reads $1-y_{s}^{2}$ De Bruyn et al. (2012), with Amhis et al. (2023b)
@@
and the average $B_{s}$ width $\Gamma_{s}$. As $y_{s}^{2}\sim 0.004$, and the relative uncertainty of $\bar{B}_{s}^{0}\rightarrow\pi^{-}D_{s}^{+}$ is an order of magnitude larger, we neglect the effect of the nonvanishing width difference for this decay.
@@
where $\delta_{X}$ is a measure of the allowed SU(3)F breaking which we set to
@@
We have altogether ten experimental data points (two branching ratios not yet being measured) and 17 real theory parameters, of which nine are magnitudes and eight are strong phases. However, we also have six constraints Eqs. (32)–(34) on these parameters, and it is indeed nontrivial if the data are in agreement with these constraints.
At the global minimum, for our null hypothesis we find $\chi^{2}=0$, i.e., a perfect fit to the data, meaning that the data can be explained with SU(3)F breaking of $\delta_{X}=30\%$, as we expect. To demonstrate the nontriviality of this result, we show in Fig. 1 the resulting $\Delta\chi^{2}$ profile as a function of $\delta_{X}$, compared to our null hypothesis.
Our fit results for $\delta_{X}=0.3$ are presented in Tables 4, 5, and 6. All errors are obtained by profiled-$\chi^{2}$ scans. Since the SU(3)F system is underconstrained, with more parameters than observables, it has a high degree of degeneracy. The parameter scans show non-Gaussianity, and appear flat around the minimum while sharply increasing near parameter boundaries. For that reason we quote parameter ranges instead of a central point with errors.
We note especially that the SU(3)F limit, which corresponds to $\delta_{X}=0$ in Fig 1 and therefore to a fit with only five parameters,
$T\,,\,\left|\frac{C}{T}\right|\,,\,\left|\frac{E}{T}\right|\,,\,\phi_{C}\,,\,\phi_{E}\,,$ (36)
results in a very high $\chi^{2}$, and is excluded. We understand this result in view of the SU(3)F limit sum rules $\varepsilon_{i}=0$, see Sec. II.2. These sum rules effectively overconstrain the fit and are clearly broken by the data (see Table 6). Therefore, the SU(3)F limit is inconsistent with the data, and SU(3)F breaking contributions have to be taken into account. Only then, does the fit have the flexibility to reach $\varepsilon_{i}\neq 0$ as required by the data. However, partial degeneracies arise concerning exactly how the required splitting of the branching ratios encoded in the $\varepsilon_{i}$ is reached. Considering, for example, the quantity $\varepsilon_{3}$ which measures the splitting between $\mathcal{B}(B^{-}\rightarrow\pi^{-}D^{0})$ and $\mathcal{B}(B^{-}\rightarrow K^{-}D^{0})$, $\varepsilon_{3}\neq 0$ can be achieved by both $T_{1}\neq 0$ or $C_{1}\neq 0$. Analogous observations hold for the other branching ratio splittings. A global fit is necessary in order to resolve these ambiguities.
@@
  • As already mentioned, the SU(3)F limit is excluded at $>5\sigma$, mainly triggered by the need for SU(3)F breaking in the tree diagrams, which is also $>5\sigma$. SU(3)F breaking in the $C$ diagrams is needed at $2.4\sigma$, while we are not yet sensitive to SU(3)F breaking in $E$ diagrams.
  • To further explore and illustrate the different importance of SU(3)F breaking in the various topologies, we consider scenarios in which SU(3)F breaking effects in only one topology at a time are taken into account, see the last three lines of Table 5. This demonstrates that SU(3)F breaking in only $C$ or $E$ does not suffice in order to obtain a good description of the data. The crucial component of the SU(3)F-breaking corrections are the ones in the tree amplitude, because only with $T_{i}\neq 0$ is the $\chi^{2}$ reduced to an acceptable level.
  • Compared to Ref. Colangelo and Ferrandes (2005), where the topological amplitudes are extracted in the SU(3)F limit, and to Ref. Chiang and Senaha (2007), where partial SU(3)F breaking effects are taken into account, we find that the general features of the results for $|C/T|$ and $|E/T|$ obtained in Colangelo and Ferrandes (2005); Chiang and Senaha (2007) are confirmed when including the first order corrections in full generality; however, the additional parameters result in larger errors for $|C/T|$.
  • The data can be described perfectly well with SU(3)F breaking of $\sim 20\%$. This conclusion holds independently of employing parameter-dependent or parameter-independent measures of SU(3)F breaking, see Fig. 1 and Table 6. Our observation agrees with the conclusions of studies of SU(3)F breaking based on factorization Bordone et al. (2020); Fleischer et al. (2011), where the factorizable SU(3)F breaking from the decay constants can be estimated as $f_{K}/f_{\pi}-1\sim 20\%$ Aoki et al. (2022), see also Ref. Jung and Mannel (2009).
  • The SU(3)F-breaking effects in $T$, $C$, and $E$ do not add up constructively in order to produce larger SU(3)F-breaking effects. This can be seen from the SU(3)F-breaking measure $\varepsilon_{4}$, which relates the $U$-spin pair $\overline{B}^{0}\to\pi^{-}D^{+}$ and $\overline{B_{s}}^{0}\to K^{-}D_{s}^{+}$, see Table 6.
@@
  • The data imply that the ratio $|E/T|$ is more strongly suppressed than $|C/T|$. However, $|C/T|$ is rather large compared to the expectation from $1/N_{c}$ counting Gronau et al. (1995). The obtained knowledge of $|E/T|$ is mainly driven by the constraint from
@@
  • The phases $\phi_{C}$ and $\phi_{E}$ are sizable, indicating large rescattering effects. The phases of the SU(3)F-breaking topologies are not yet very well known.
  • Because of lack of data we cannot yet test the isospin relation $s_{1}=1$; see Eqs. (23, 29). From our fit we obtain the corresponding yet unmeasured branching ratios as
@@
    The interfering $B_{s}$ decays into the same final states are in this case of the same order in Wolfenstein-$\lambda$, as it is also the case for $\overline{B}_{s}^{0}\rightarrow K^{\mp}D_{s}^{\pm}$ decays, see above. Estimating the impact from $\Delta\Gamma_{s}\neq 0$ as $\sim 20\%$, as suggested by the size of the effect for $\mathcal{B}(\overline{B}_{s}^{0}\rightarrow K^{-}D_{s}^{+})$, we predict
@@
    Future measurements of $\mathcal{B}(\overline{B}^{0}_{s}\to\pi^{-}D^{+})$ and $\mathcal{B}(\overline{B}^{0}_{s}\to\pi^{0}D^{0})$ would allow one to test their isospin symmetry relation and at the same time improve the bounds on $|E/T|$.